# Random Schreier graphs as expanders

### v6

@@ 4 Graphs from Toeplitz matrices
We now turn to modify our construction in order to reduce further the number of used random bits. Instead of taking random invertible matrices, we take Toeplitz matrices. Recall that a matrix $T=(t_{ij})$ is called a Toeplitz matrix if $t_{ij}$ depends only on the difference $j-i$, i.e., if
$T=\begin{pmatrix}a_{0}&a_{1}&a_{2}&\cdots&\cdots&\cdots&a_{k-1}\\ a_{-1}&a_{0}&a_{1}&\ddots&&&\vdots\\ a_{-2}&a_{-1}&a_{0}&\ddots&\ddots&&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \vdots&&\ddots&\ddots&a_{0}&a_{1}&a_{2}\\ \vdots&&&\ddots&a_{-1}&a_{0}&a_{1}\\ a_{-k+1}&\cdots&\cdots&\cdots&a_{-2}&a_{-1}&a_{0}\\ \end{pmatrix}$
for some $a_{-k+1},\ldots,a_{0},\ldots,a_{k-1}$.
If we choose the elements of such a matrix randomly from a field of characteristics $q$, then we get an invertible matrix with probability $1-1/q$ (see [14]), so it is even easier to generate them. A Toeplitz matrix is defined by $12k-$ 1 elements of the field (instead of $k^{2}$ for a general matrix), so this is also an advantage. (One more technical advantage is that we can multiply a vector by a Toeplitz matrix very fast, since it is essentially the convolution.)
Figure 5: Empirical distributions of the second largest eigenvalue of 5000 graphs obtained from Toeplitz matrices (labeled “TP’) (for $\mathbb{F}_{2}^{14}$ and $\mathbb{F}_{7}^{5}$) compared to the corresponding empirical distributions for random linear permutations in the same vector spaces (labeled “GL”).
In Figure 5 we return to the case of undirected regular graphs, and compare random graphs obtained using $15$ random matrices (either random invertible matrices or random invertible Toeplitz matrices). Two upper curves (yellow and blue) use $\mathbb{F}_{2}^{14}$; one can see that Toeplitz matrices give bigger (and less concentrated) eigenvalues. The same can be seen from two bottom curves (red and green); we see also that the decreasing the dimension and increasing field size increase significantly the variance (but this is not the case for the mean in our example).
@@ 5 Computer-assisted theoretical bounds
The numerical experiments show that the constructions discussed in the previous sections allow to sample good spectral expanders of reasonably small size (e.g., with $n\sim 10^{4}$ vertices, when we can keep the entire graph in the memory of a computer). This might be enough for practical applications: we sample at random a multiset $S$ of matrices, produce the corresponding graph, and compute its eigenvalues. If we obtain a good spectral expander, we can fix the used random bits and plug this instance of a graph in our application.
@@
However, in some applications we may need “strongly explicit” expanders with much larger number of vertices (e.g., with $n\sim 2^{200}$). In this case we cannot produce the entire matrix of the graph, although we still can keep in the memory an index of one vertex and we can compute efficiently the indices of its neighbours. In this case, we are not able to compute numerically the spectrum of the graph. So it would be helpful to have a theoretical result saying that most of (pseudo)random graphs (for some model of sampling) are good spectral expanders. Another reason why we are interested in theoretical bounds (even for graphs of small size) is less practical: we would like to explain the phenomena that we observe in numerical experiments.
In what follows we adapt to our setting the proof technique from [6] (we use the trace method and analyse the probability of getting a closed path during a random walk). We use it to prove bounds on the expected second largest eigenvalue of the adjacency matrix of a sampled graph. Our proofs involve heavy computations, so they are in some sense computer-assisted. All our theoretical results (except Theorem 3) only apply to $\mathbb{F}_{2}$ (though the experimental results show that the construction works well for bigger fields, see Figures 1 and 3).
5.1 Spectral bound for regular graphs
@@
Theorem 1.
Let $G$ be a Schreier graph of $GL_{k}(\mathbb{F}_{2})$ acting on $(\mathbb{F}_{2}^{k})^{*}$ with the random multiset $S$, $n=2^{k}-1$ (with $k\geq 2$), and degree $2d=2|S|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalized adjacency matrix. Consider the following recurrent relation:
@@
$\mathbf{E}(|\mu_{2}|)\leq\\ \bigg{(}\bigg{(}\frac{1}{2d}\bigg{)}^{2m}n\bigg{[}\sum_{i=1}^{m}\binom{d}{i}\frac{(2m)!}{(2m-2i)!}\bigg{[}(2^{i}-1)(x_{3}(i)+x_{4}(i))\frac{2}{n}\\ +\bigg{(}(x_{3}(i)+x_{4}(i))\frac{1}{n}+\frac{2^{i}}{(i+1)!}\bigg{(}x_{3}(i)\frac{5}{n}+x_{4}(i)\bigg{)}\bigg{)}\bigg{]}\\ +X_{1}(1,2m,d)\frac{1}{n}+x_{3}(0)\frac{5}{n}+x_{4}(0)\bigg{]}-1\bigg{)}^{\frac{1}{2m}}.$ (1)
The bound (1) looks messy, and the asymptotic analysis may be difficult. However, this formula becomes useful when we need to calculate (with the help of a computer) a bound for a specific graph of relatively small size (e.g., less than $2^{200}$ vertices).
To apply this theorem for given $k$ and $d$, we choose the value of $m$ such that (1) implies the strongest bound for $\mathbf{E}(|\mu_{2}|)$. We discovered that for several “reasonable” values of $k$ and $d$, with $2m$ chosen close to $k$, this formula implies a non trivial bound for $\mu_{2}$.
5.2 Spectral bound for bipartite graphs
Theorem 2.
Let $G$ be a Schreier bipartite graph of $GL_{k}(\mathbb{F}_{2})$ ($k\geq 2$) acting on $(\mathbb{F}_{2}^{k})^{*}$ with the random multiset $D$, $|D|=d$, and $n=|(\mathbb{F}_{2}^{k})^{*}|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalized adjacency matrix. Consider the relation
$Y_{p}(c,\ell,d)=\left\{\begin{array}[]{ll}1\mbox{ if }c=0\mbox{ and }\ell=0;\\ 0\mbox{ if }p>\ell;\\ \sum_{i=c}^{\lfloor\frac{\ell}{p}\rfloor}\binom{d}{i}(p!)^{-i}\frac{\ell!}{(\ell-ip)!}Y_{p+1}(0,\ell-ip,d-i)\mbox{ otherwise.}\end{array}\right.$
@@ 5.2 Spectral bound for bipartite graphs
Again, the expression in (2) is difficult to analyse, but it helps to compute specific (not asymptotic) bounds for some specific values of parameters.
@@
5.3 A simple asymptotic bound
In this work we do not focus on asymptotic behaviour of large graphs. However, we observe that at least in some setting, the used technique easily gives some asymptotic bounds. This bound applies to the construction of a Schreier graph $GL_{k}(\mathbb{F}_{q})$ with any dimension $k$ (including $k=1$) and with any prime $q$, which is not the case of the previous ones (in Theorem 1 and Theorem 2 we use very substantially the fact that the group is not abelian, which is false in dimension one). The bound applies to graphs with a pretty large degree, $d=\Omega(\ln n)$ (see next section for an estimate of the constant hidden in the $\Omega$ notation).
Theorem 3.
Let $G=Sch(GL_{k}(\mathbb{F}_{q})\circlearrowleft(\mathbb{F}_{q}^{k})^{*},S)$ graph of size $n=q^{k}-1$ and whose degree is $2d=2|S|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalised adjacency matrix. Then,
$\mathbf{E}(|\mu_{2}|)\leq e\sqrt{\frac{\ln n}{d}}$
where $\ln$ is the natural logarithm and $e$ its base.
5.4 Computer-assisted upper bounds for the second eigenvalue
Our experimental observations show that the second eigenvalue of a random Schreier graph from our construction is very close to the theoretical optimum (among all regular graphs). This phenomenon remains unexplained. What is even more frustrating, and we cannot guarantee that (pseudo) random graphs from similar families but with a large number of vertices (e.g., with $n\sim 2^{200}$) have good spectral properties. However, Theorem 1 and Theorem 2 imply some nontrivial bounds for the second eigenvalues of those graphs. In Table 1 we show several examples of such bounds and compare them with the Ramanujan’s threshold.
We observed that the bound from Theorem 1 becomes non-trivial (i.e., the bound for the normalised second eigenvalue $\mu_{2}$ is less than $1$) for $d=\Omega(k)$ ($d$ is the degree of the graph). The experiments suggest that for $d=2k$ and large number of vertices this theorem gives a bound that converges to a constant around $0.7$ (although we do not have a formal justification of this hypothesis). This bound gives more interesting implications for larger degree of the graph (compared with parameter $k$, e.g., for $d\geq k^{2}$). If we compare the bounds from Theorem 1 and 2 in terms of number of random matrices used, the first bound could seem slightly better than the second (though pretty similar). However, if we compare them in terms of degree, Theorem 2 gives a better bound, see Table 1. This is because increasing the number of matrices used improves the bound in both theorems and we need twice as much random matrices for bipartite graphs to get the same degree.
$(k,d)$ Th.1 ratio Th.2 ratio
$(14,30)$ 0.7607 2.1190 0.5787 1.6122
$(14,60)$ 0.5033 1.9657 0.3923 1.5325
$(14,120)$ 0.3389 1.8640 0.2687 1.4781
$(20,40)$ 0.7758 2.4847 0.5741 1.8387
$(25,100)$ 0.4969 2.4971 0.3718 1.8685
$(30,1000)$ 0.1435 2.2709 0.1128 1.7846
$(40,1000)$ 0.1618 2.5604 0.1251 1.9792
$(60,120)$ 0.7377 4.0577 0.5086 2.7977
$(200,400)$ 0.7016 7.0251 0.4592 4.5977
@@ 5.4 Computer-assisted upper bounds for the second eigenvalue
Table 1: This table shows the upper bound for the expected value of $\mu_{2}$ computed with Theorems 12 for a few examples of parameters values (the dimension $k$ and the degree $d$). Note that the regular non bipartite graphs are constructed with $d/2$ random matrices whereas the bipartite ones used $d$ random matrices. We display also the ratio between the Ramanujan’s threshold and the values given by the bounds ($\mu_{2}/\frac{2\sqrt{d-1}}{d}$). The first three lines correspondent to graphs that we could test in our experiments
6 Final comments
In this paper we study a pseudo-random construction of spectral expanders represented as Schreier graphs. The experimental results suggest that these graphs have nearly optimal value for the second largest eigenvalues (close to $2\sqrt{d-1}$ for regular graphs of degree $d$ and, respectively, to $\sqrt{d_{1}-1}+\sqrt{d_{2}-1}$ for bi-uniform bipartite graphs with degrees $d_{1}$ and $d_{2}$), and this is true not asymptotically but for graphs of reasonably small sizes relevant for practical applications.
A possible direction of the future research is a reduction of the number of random bits used to produce each graph. It would be interesting to reduce the number of these bits from $O(d\log^{2}n)$ to $O(d\log n)$ and improve computational complexity of the construction. Our experiments show that a promising direction is a construction based on Toeplitz matrices.
The “computer assisted theoretical bounds” that we provide are a small step to an explanation of the obtained experimental results. Instead of more traditional asymptotic bounds, we focused on theoretical bounds that can be calculated (possibly with the help of computer) for graphs with specific values of parameters, including the values of parameters that appear in our numerical experiments. In the appendix (A.4), we discuss limitations of the used method. It seems that a better theoretical explanation of the behaviour of random Schreier graphs of $GL_{k}(\mathbb{F}_{q})$ will require a subtler analysis.
We would also like to draw attention to the observation that useful theoretical bounds and estimates can be obtained with the help of “hideous” formulas combined with computer computations. This approach contradicts the spirit of classical mathematics, where we appreciate proofs that are elegant and meaningful for a human. However, these techniques can be useful to justify properties of objects (e.g., of pseudo-random graphs) with specific parameters, which may be necessary in practical applications.
Acknowledgements
The author thanks his PhD advisers, Andrei Romashchenko and Alexander Shen, for providing several ideas, useful remarks and help on this work.
Appendix A Proofs of the main results
A.1 Proofs of Theorems 1 and 3

### v7

@@ 4 Graphs from Toeplitz matrices
We now modify our construction in order to reduce further the number of random bits. Instead of taking random invertible matrices, we take invertible Toeplitz matrices. Recall that a matrix $T=(t_{ij})$ is called a Toeplitz matrix if $t_{ij}$ depends only on the difference $j-i$.
If we choose the elements of such a matrix randomly from a field of prime size $q$, then we get an invertible matrix with probability $1-1/q$ (see [14]), so it is even easier to generate them. A Toeplitz matrix is defined by $12k-$ 1 elements of the field (instead of $k^{2}$ for a general matrix), so this is also an advantage. (One more technical advantage is that we can multiply a vector by a Toeplitz matrix fast since it is essentially the convolution.)
Figure 5: On the left: distributions of second largest eigenvalues of 5000 graphs obtained from Toeplitz matrices (TP) for $\mathbb{F}_{2}^{14}$ and $\mathbb{F}_{7}^{5}$ compared with the corresponding empirical distributions for random linear permutations in the same vector spaces (GL). On the right, all graphs are from Toeplitz matrices over $\mathbb{F}_{2}$.
In Figure 5 (left) we return to the case of undirected regular graphs, and compare random graphs obtained using $15$ random matrices (either random invertible matrices or random invertible Toeplitz matrices). Two upper curves (yellow and blue) use $\mathbb{F}_{2}^{14}$; one can see that Toeplitz matrices give bigger (and less concentrated) eigenvalues. The same can be seen from two bottom curves (red and green); we see also that decreasing the dimension and increasing the field size increase significantly the variance (but not the mean, see below). The difference between Toeplitz and general invertible matrices is visible but small. Hence, Topelitz matrices are suitable for producing good spectral expanders (at least with the parameters of the experiments). On the right of Figure 5, we observe that the parity of the dimension of Toeplitz matrices used affects the mean of the distribution (the variance behaves as expected). Odd dimensions tend to produce better spectral expanders (this can be also observed on the left part of the Figure, red curve). We could not find any explanation of this phenomenon.
@@ 5 Computer-assisted theoretical bounds
The numerical experiments show that the constructions discussed in the previous sections allow to sample good spectral expanders of reasonable size (say, with $n\sim 10^{4}$ vertices) where we can compute the second eigenvalue. However, in some applications we may need “strongly explicit” expanders with much larger number of vertices (see e.g. [9]). In this case we cannot produce the entire matrix of the graph (nor compute the eigenvalues), although we still can keep in the memory an index of one vertex and compute efficiently the indices of its neighbors.
It is thus useful to have theoretical guaranties about the eigenvalues when we cannot compute them. To do so, one can use the trace method [6] for the case of Cayley (or arbitrary Schreier) graphs, following [3]. This gives the following bound (see also [22] and [8]):
Proposition 1
Let $G$ be a Schreier graph constructed using $d$ elements of a group acting on an $n$-element set (so it has $n$ vertices and degree $2d$). Let $\mu_{2}$ be the second largest eigenvalue of its normalized adjacency matrix. Then
$\mathbf{E}(|\mu_{2}|)\leq e\sqrt{\frac{\ln n}{d}}$
where $\ln$ is the natural logarithm and $e$ its base. The same bound holds for graphs constructed with $d$ randomly chosen invertible Toeplitz matrices. The statement is also valid for bipartite graphs of $2n$ elements of degree $d$ (produced with $d$ random elements).
Proof sketch. The proof of this bound goes as follows. To get an upper bound for the eigenvalues of some symmetric real matrix $M$, we can use upper bounds for the trace of $M^{m}$ for some even $m$ (that will be chosen later). Indeed, the trace is the sum of the $m$th powers of the eigenvalues. In our case $M$ is the (normalized) adjacency matrix, and we know the first eigenvalue $1$. Hence, given an upper bound for the trace of $M^{m}$ we get some bound for the second eigenvalue. (In fact, we will get a bound for expected value of $\mu_{2}^{m}$, but this will imply a bound for the expected value of $\mu_{2}$ due to convexity.)
Trace of $M^{m}$ for a normalized adjacency matrix $M$ has a natural interpretation in terms of random walks: the $i$th diagonal element of the matrix is the probability to return to $i$ after $m$ steps of a random walk (starting from $i$). We need a bound for the expected value of this trace (for a random choice of $d$ elements of the group). So we have two independent choices: (a) choice of $d$ group elements $h_{1},\ldots,h_{d}$; (b) choice which of $h_{1},\ldots,h_{d},h_{1}^{-1},\ldots,h_{d}^{-1}$ will be used along the walk ($m$ choices from $2d$-element set). We can exchange the order of averaging and first take a random word of length $m$ over the alphabet with $2m$ symbols $h_{1},\ldots,h_{d},h_{1}^{-1},\ldots,h_{d}^{-1}$, and only after that choose the random values (group elements) for $h_{1},\ldots,h_{d}$.
For example, with some probability we get after the first stage a word of length $m$ where the letters cancel each other completely. Then on the second stage we remain in the same vertex with probability $1$. But if $m$ is significantly smaller than $d$, then with high probability we get a non-canceling word, and, moreover, a word where some letter $h_{i}$ appears only once (as $h_{i}$ or $h_{i}^{-1}$). In this case the second stage probability to return back is $1/n$ (since this unique application of $h_{i}$ will make the distribution uniform, due to transitivity of the action). If we combine these remarks carefully (and take $m\approx\ln n$), we get the required bound. See Appendix 0.B.1 for details.
A more subtle classification of the words of length $m$ over the alphabet composed of the symbols $\{h_{1},\ldots,h_{d},h_{1}^{-1},\ldots,h_{d}^{-1}\}$ allows us to improve this bound (Theorem 0.A.1) for Schreier graphs of $\textit{GL}_{k}(\mathbb{F}_{2})$. Unfortunately, the improved bound is quite complicated (and even its asymptotic behavior is not clear). Both the bound and its proof are included in the Appendix. However, using computer assistance, we can compute this improved bound for specific values of parameters ($k$ and $d$).
$(k,d)$ Improved Proposition 1 Ramanujan
$(14,15)$ 0.7607 2.1863 0.3590
$(14,30)$ 0.5033 1.5459 0.2560
$(14,60)$ 0.3389 1.0931 0.1818
$(20,20)$ 0.7758 2.2631 0.3122
$(25,50)$ 0.4969 1.6002 0.1989
$(30,500)$ 0.1435 0.5543 0.0632
$(40,500)$ 0.1618 0.6401 0.0632
$(50,1000)$ 0.1217 0.5060 0.0447
$(60,60)$ 0.7377 2.2631 0.1818
$(200,200)$ 0.7016 2.2631 0.0998
Table 1: Upper bounds for the expected value of $\mu_{2}$ provided by Proposition 1 and its improved version (see Appendix) for a few choices of parameters. We display also the Ramanujan’s threshold for comparison. The first three lines correspond to graphs that we tested in our experiments.
Looking at this table, we see that the improved bound makes sense in several cases where Proposition 1 does not give anything (the bound is greater than the trivial bound $1$). However, even the improved bound is far from being tight (recall that our experiments, when feasible, exhibit a rather small difference between the second eigenvalue and Ramanujan’s threshold). Moreover, we can observe some convergence behavior when $d=k$. This would imply that the asymptotic behavior of the improved bound is similar to that of Proposition 1. Similar bounds can be obtained for the bipartite case (see Appendix), and these observations remain valid.
6 Final comments
In this paper we study a pseudo-random construction of spectral expanders using Schreier graphs. The experimental results suggest that these graphs provide almost optimal value for the second largest eigenvalue (close to $2\sqrt{d-1}$ for regular graphs of degree $d$ and to $\sqrt{d_{1}-1}+\sqrt{d_{2}-1}$ for bi-uniform bipartite graphs with degrees $d_{1}$ and $d_{2}$), and are potentially useful for the applications.
Still there is no theoretical explanation for this phenomenon: the existing bounds are quite weak, even in the improved version (and these improvements do not give a closed answer, just a recurrent formula that allows us to compute the bound).
The possible reduction of the number of random bits used and the computational complexity via Toeplitz matrices looks promising (as the experiments show) but a comprehensive theoretical analysis is missing. The only proven bound that applies in the setting with Toeplitz matrices is Proposition 1, which gives nontrivial bounds only if the degree of the graph is bigger than logarithm of the number of vertices.
Appendix 0.A Theoretical bounds
In appendix we discuss computer assisted theoretical bounds that we prove for the eigenvalues of random graphs that we observed in numerical experiments. These bounds provide a small step to a theoretical explanation of the obtained experimental results. Instead of more traditional asymptotic bounds (as in [11]), we focused on the bounds that can be calculated (possibly with the help of computer) for graphs with specific values of parameters, including the values of parameters that appear in our numerical experiments.
We would also like to draw attention to the observation that useful theoretical bounds and estimates can be obtained with the help of “hideous” formulas (see Theorem 0.A.1 and Theorem 0.A.2 below) combined with computer computations. This approach contradicts the spirit of classical mathematics, where we appreciate proofs that are elegant and meaningful for a human. However, this technique can be useful to justify properties of objects (e.g., of pseudo-random graphs) with specific parameters, which may be necessary in practical applications.
0.A.1 Spectral bound for regular graphs
@@
Theorem 0.A.1
Let $G$ be a Schreier graph of $\textit{GL}_{k}(\mathbb{F}_{2})$ acting on $(\mathbb{F}_{2}^{k})^{*}$ with the random multiset $S$, $n=2^{k}-1$ (with $k\geq 2$), and degree $2d=2|S|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalized adjacency matrix. Consider the following recurrent relation:
$X_{p}(c,l,d)=\left\{\begin{array}[]{ll}1\mbox{ if }c=0\mbox{ and }l=0;\\ 0\mbox{ if }p>l;\\ \sum_{i=c}^{\lfloor\frac{l}{p}\rfloor}\binom{d}{i}\bigg{(}\frac{2^{p}}{p!}\bigg{)}^{i}\frac{l!}{(l-pi)!}X_{p+1}(0,l-pi,d-i)\mbox{ otherwise.}\end{array}\right.$
@@
$\mathbf{E}(|\mu_{2}|)\leq\bigg{(}\bigg{(}\frac{1}{2d}\bigg{)}^{2m}n\bigg{[}\sum_{i=1}^{m}\binom{d}{i}\frac{(2m)!}{(2m-2i)!}\bigg{[}(2^{i}-1)(x_{3}(i)+x_{4}(i))\frac{2}{n}\\ +\bigg{(}(x_{3}(i)+x_{4}(i))\frac{1}{n}+\frac{2^{i}}{(i+1)!}\bigg{(}x_{3}(i)\frac{5}{n}+x_{4}(i)\bigg{)}\bigg{)}\bigg{]}\\ +X_{1}(1,2m,d)\frac{1}{n}+x_{3}(0)\frac{5}{n}+x_{4}(0)\bigg{]}-1\bigg{)}^{\frac{1}{2m}}.$ (1)
The bound (1) looks messy, and the asymptotic analysis may be difficult. However, this formula becomes useful when we need to calculate (with the help of a computer) a bound for a specific graph of relatively small size. Let us mention that a “relatively small” graph is not necessary a graph that can be stored in computer’s memory. In practice, we may need graphs with the number of vertices exponentially larger than the available memory (e.g., we may need an expander graph with $2^{256}$ vertices, where every vertex of the graph represents a string of $32$ bytes) [^3] One the well-known application of expanders is a deterministic error amplification, which helps to reduce the error probability in a randomized algorithm without increasing the number of random bits, see [13, Section 1.1.3]. In applications of this type, one would need very large expander graphs, so that we cannot store the entire graph in the memory of a computer and only keep in mind the indices of a few vertices. In this case, we would need a strongly explicit construction of an expander: given an index of a vertex, we should be able to compute efficiently the list of its neighbors without storing the whole graph. Our constructions are strongly explicit, and they, probably, can be used for applications of this kind. However, we cannot compute numerically the eigenvalues of graphs whose size is exponential compared with the memory of a computer. So in this case we need theoretical bounds for the spectrum of sampled graphs..
To compute this bound for given $k$ and $d$, we choose the value of the size of the walk ($2m$) such that (1) implies the strongest bound for $\mathbf{E}(|\mu_{2}|)$. The exact value of $m$ that minimises the bound is unknown, but we suspect it to be logarithmic as in Proposition 1. In order to compute this bound for some $k$ and $d$, we set $m=1$ and compute (1) right hand side several times, incrementing $m$ until it ceases decreasing (1) (the bound will converge to $1$ after reaching its minimum when $m$ grows). In all our computations, this happens for $m\leq k$. We discovered that for several “reasonable” values of $k$ and $d$, this formula implies a non trivial bound for $\mu_{2}$ (see Table 1).
0.A.2 Spectral bound for bipartite graphs
Theorem 0.A.2
Let $G$ be a Schreier bipartite graph of $\textit{GL}_{k}(\mathbb{F}_{2})$ ($k\geq 2$) acting on $(\mathbb{F}_{2}^{k})^{*}$ with the random multiset $D$, $|D|=d$, and $n=|(\mathbb{F}_{2}^{k})^{*}|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalized adjacency matrix. Consider the relation
$Y_{p}(c,l,d)=\left\{\begin{array}[]{ll}1\mbox{ if }c=0\mbox{ and }l=0;\\ 0\mbox{ if }p>l;\\ \sum_{i=c}^{\lfloor\frac{l}{p}\rfloor}\binom{d}{i}(p!)^{-i}\frac{l!}{(l-pi)!}Y_{p+1}(0,l-pi,d-i)\mbox{ otherwise.}\end{array}\right.$
@@ 0.A.2 Spectral bound for bipartite graphs
Again, the expression in (2) is difficult to analyse, but it helps to compute specific (not asymptotic) bounds for some specific values of parameters. In Table 2, we show the values given by the bound for some arbitrary parameters. The bounds are computed the same way as for regular non bipartite graphs. The bound for bipartite graph is stronger than that of Theorem 0.A.1. Indeed in order to achieve the same degree, twice more random matrices are used.
@@
$(k,d)$ Th.0.A.2 Ramanujan
$(14,30)$ 0.5787 0.3590
$(14,60)$ 0.3923 0.2560
$(14,120)$ 0.2687 0.1818
$(20,40)$ 0.5741 0.3122
$(25,100)$ 0.3718 0.1989
$(30,1000)$ 0.1128 0.0632
$(40,1000)$ 0.1251 0.0632
$(60,120)$ 0.5086 0.1818
$(200,400)$ 0.4592 0.0998
@@
Table 2: This table shows the upper bound for the expected value of $\mu_{2}$ computed with Theorem 0.A.2 for a few examples of parameters values (the dimension $k$ and the degree $d$). Note that the bipartite are constructed by sampling at random $d$ matrices. We display also the Ramanujan’s threshold($\frac{2\sqrt{d-1}}{d}$). The first three lines correspondent to graphs that we could test in our experiments
Appendix 0.B Proofs
0.B.1 Proof of Proposition 1